what is the mass of a sample of guanine C5H5N5O in which the total number of atoms is 7.43 X 1023
Answer #1

The compound is Guanine = C5H5N5O

molar mass of Guanine = C5H5N5O = 151.13 gram/mol

one mole of compound contains Avagadro number of atoms. i,e 6.023 x10^23 atoms

Total number of atoms present in C5H5N5O = (5+5+5+1)X 6.023X10^23 = 16 X 6.023X10^23 atoms

mass of one mole of compound = 151.13 grams

mass of 16 x 6,023x10^23 atoms of C5H5N5O = 151.13 grams

mass of 7.43 x10^23 atoms of C5H5N5O = ?

                                                                 = 7.43 x10^23 x 151.13 / 16x6.023x10^23 = 70.18 grams

mass of 7.43 x10^23 atoms of C5H5N5O = 70.18 grams.